OpenStudy (logo):

Help pls will award medal! Simplify the expression -5 + i/ 2i

OpenStudy (logo):

Does it work out like this? -5+i/2i -2i*-5+i/2i*-2i 10i - 2i^2/-4i^2 10i-2/4 -0.5 + 10i?

OpenStudy (karleighxxvictoria):

-5+i square/2

OpenStudy (karleighxxvictoria):

well its really 10i-2 (divided) 4

OpenStudy (logo):

so 10i - 2/4 is the answer? so my steps are correct?

OpenStudy (karleighxxvictoria):

yes

OpenStudy (tkhunny):

First, you have to be clear what it is you are writing. You wrote this: -5 + i/ 2i which means \(-5 + \dfrac{i}{2i} = -5 + 1/2 = -9/2\) I'm guessing you meant : (-5 + i)/ 2i which means \(\dfrac{-5+i}{2i} = \dfrac{(-5+i)}{2i}\cdot\dfrac{i}{i} = \) \(= \dfrac{(-5i + i^{2})}{2i^{2}} = \dfrac{(-5i + (-1))}{2(-1)} = \dfrac{-5i - 1}{-2} = \dfrac{5i + 1}{2}\) You can continue to meet odd requirements from your teacher, after that, maybe 1/2 + (5/2)i. Second, you should simplify before you think you are done.

OpenStudy (tkhunny):

Note: It isn't necessary to use the Complex Conjugate unless there is a Real Portion. 1 + i requires multiplication by 1- i. 0 + i does not require multiplication by 0 - i. Simply "i" will do.

OpenStudy (logo):

@karleighxxvictoria and @tkhunny thx! I got it!